OpenStudy (soldierbxrnes):

Add the polynomials. (82x8+21x2-6)+(18x+7x8-42x2+3)

Directrix (directrix):

What is the sum of: 82x^8 + 7x^8 ?

OpenStudy (soldierbxrnes):

90x^8??

Directrix (directrix):

82x^8 + 7x^8 = 89 x^8 Now, do this: 21 x^2 - 42 x^2 = ? And, then this: -6 + 3 =

Directrix (directrix):

21 x^2 - 42 x^2 = ?

OpenStudy (soldierbxrnes):

-21x^2

Directrix (directrix):

Correct

Directrix (directrix):

: -6 + 3 =

OpenStudy (soldierbxrnes):

-3

Directrix (directrix):

Correct. Before we put these together, a question about the problem. Is there supposed to be an x term in the first expression: (82x8+21x2-6) The second part has an x term so I am wondering. There does not have to be an x term in the first expression but please check.

OpenStudy (soldierbxrnes):

no there isn't an x term

Directrix (directrix):

Okay. Putting the like terms together: 82x8 + 7x8 +21x2 - 42x2 + 18x -6 + 3 = 89 x^8 - 21x^2 + 18x - 3 Final Answer - Do you agree?

OpenStudy (soldierbxrnes):

yes i got the exact same answer thank you so much!

Directrix (directrix):

You are welcome. Note that there is no like term to add to the 18x so it remains "solo."

OpenStudy (soldierbxrnes):

Quick question, what do i do when the terms are like this: x^9? does that mean its basically 1x^9?

OpenStudy (skullpatrol):

Yes.

OpenStudy (skullpatrol):

You are using the distributive property to add similar terms. For example $$ x^9 + 2x^9 = (1+2)x^9 = 3x^9$$

OpenStudy (skullpatrol):

Good question :-)

OpenStudy (skullpatrol):

To add (or subtract) polynomials you add (or subtract) their similar terms. Similar terms are monomials that are exactly alike or that differ only in their numerical coefficients.